# A Federated Distributionally Robust Support Vector Machine with Mixture of Wasserstein Balls Ambiguity Set for Distributed Fault Diagnosis

### v1

@@ 5.3.2 ADMM Algorithm
We begin by studying the complexity of solving the strongly convex variant of the local model problem in (12d) solved by each client during each iteration $t$. Using the $\ell_{1}$-norm, we can express our problem as a QCQP with $N_{g}+2P+3$ decision variables (including slack variables) and $32N_{g}+2P+$ 3 constraints. This can be solved using the barrier method equipped with the log barrier function and Newton updates. For this class of problems, the aforementioned method requires $\mathcal{O}(\sqrt{C}\log(\epsilon_{2}^{-1}))$ iterations to reach an $\epsilon_{2}$-solution, where $C$ is the number of constraints [43]. Each iteration has an arithmetic complexity of $\mathcal{O}([C+D]D^{2})$, where $D$ is the number of decision variables [43]. Therefore, solving the problem via the previously discussed method has a complexity of
$\mathcal{O}([2N_{g}+2P+3]^{0.5}[3N_{g}+4P+6][N_{g}+2P+3]^{2}\log(\epsilon_{2}^{-1})).$
By eliminating scalar multipliers and constants, the problem can be simplified to the following:
@@ 6.1.1 Global Parameters
We make multiple observations from the experimental results. Firstly, we observe that the SM algorithm outperforms all the other distributed models, and achieves very similar performance to the central model. Moreover, it is less sensitive to changes in $\gamma$ than the ADMM algorithms are to changes in $\rho$ as it maintains its peak performance for a wider range of the parameter. This performance, however, comes at the cost of training time, as it requires more iterations than the ADMM algorithms to reach its peak performance in many cases. Moreover, the performance of the SM algorithm seems to be much less affected by client and class imbalance and their combination than ADMM algorithms, making it a good choice in settings where such imbalance is known to exist. Secondly, we make the surprising observation that the FDR-SVM trained via the ADMM algorithm is outperformed by its regularized counterpart (ADMM-REG) in all but the noisy labels setting. However, this can be explained by noting that the key difference between the two models is that the DR model accounts for label uncertainty via the label flipping cost parameter $\kappa$, unlike the regularized model [7]. This can make the FDR-SVM equipped with a relatively low $\kappa$ too conservative where all the training labels are correct. Indeed, this is further confirmed when we note that the FDR-SVM slightly outperforms its regularized counterpart in the noisy label setting. We note, however, that with appropriate local parameter choices ADMM can still outperform ADMM-REG as we show in the local parameter experiment. Additionally, we observe that ADMM-SC is consistently outperformed by ADMM and ADMM-REG, which can be expected due to the redundant regularization provided by the additional strongly convex term. Indeed, observe that the performance of ADMM-SC drops more significantly than that of ADMM and ADMM-REG as $\rho$ increases. This is because increasing $\rho$ leads to increasing $\tau$, which increases the magnitude of redundant regularization and impacts model performance. Finally, and also very surprisingly, we observe that ADMM and ADMM-REG very slightly outperform the centralized model in the noisy label setting. This is unexpected as the centralized model typically outperform distributed ones. This may indicate that the central model is highly sensitive to its parameters in this setting. Alternatively, we could explain this result by noting that our proposed FDR-SVM uses a different ambiguity set from the central model’s. This may suggest that our proposed MoWB ambiguity set is more successful at modeling uncertainty than the traditional Wasserstein ball in certain settings.
@@ 6.1.2 Local Parameters
A key observation that can be made from the results is that with low $\kappa$ values, ADMM is outperformed by ADMM-REG. However, ADMM matches, and in some cases, exceeds the performance of ADMM-REG as $\kappa$ increases. This confirms our hypothesis that ADMM was outperformed by ADMM-REG in the global parameter experiment due to the over-conservatism of the DR model equipped with low $\kappa$ values. This result indeed confirms that the FDR-SVM will always perform just as well as its regularized counterpart, if not better, in many settings including uncertain training labels. We also observe that SM is less sensitive than ADMM to the choice of $\kappa$ for all settings except for the noisy label setting, exhibiting almost opposite behavior to ADMM. Moreover, SM is also relatively highly sensitive to the choice of local Wasserstein radius. Indeed, in most settings the performance of SM improves as the local radius decreases. However, for noisy labels performance drops as the local radius decreases. This is likely due to the need for larger local ambiguity sets to account for both feature and label uncertainty. Additionally, we observe again that SM’s peak performance is higher than that of ADMM in all settings except for noisy labels. This is because the results of the global parameter experiment show that SM achieves peak performance in this setting with $\gamma=1\times 10^{1}$. However, in this experiment we test with $\gamma=1\times 10^{2}$. Thus, by referencing the previous experiment’s results we see that SM would still perform just as well as ADMM given appropriately chosen parameters. Another observation we make is that ADMM-SC is once again outperformed by all the other models for most parameter values due to redundant regularization. A surprising result, however, is that in some settings the error bars of the ADMM-SC exceed all the other models in performance, even the central model. This again suggests that the MoWB ambiguity set may be able to characterize the uncertainty better than the traditional Wasserstein ball in some settings.
@@ 6.2 Experiment 2: Scalability Experiment
This experiment examines the scalability of the algorithms and models under study using simulation data generated via the make_classification module [51]. The data distribution across clients and across classes is identical to that utilized in the nominal setting of the previous experiment. We test the following settings: a) Increasing clients [fixed training samples]: $N=1000$, $P=4$, $G\in\{10,20,30,40,50\}$; b) Increasing clients [increasing training samples]: $N=100G$, $P=4$, $G\in\{10,20,30,40,50\}$; c) Increasing training samples: $G=10$, $P=4$, $N\in\{1000,1500,2000,2500,3000\}$, and d) Increasing features: $N=4$, $G=10$, $P\in\{4,6,8,10,12\}$. It is important to note that the subgradient method lacks a practically implementable stopping criterion [59], while no stopping criterion is provided for multi-block ADMM in [49]. Thus, we opt for evaluating scalability in the same fashion for all algorithms: we compute the mean runtime in seconds over $50$ repetitions required by each algorithm to converge to the model with the highest mCCR. During each repetition, we use new randomly generated training and test sets. For the SM algorithm we test performance for $T\in\{140,180,220\}$ and $\gamma\in\{1\times 10^{b}\}_{b=1}^{3}$. Meanwhile, for the ADMM algorithms we test performance for $T\in\{10,20,30\}$ and $\gamma\in\{1\times 10^{b}\}_{b=-3}^{-1}$. For all algorithms, we fix $\varepsilon_{g}=\frac{1}{10N_{g}}$ and $\kappa=0.25$. The central model we trained is tuned in an identical fashion to that from the previous experiment, and the reported runtime is that taken to solve the optimization problem. The results of this study are reported in Figure 3.
@@ 6.3 Experiment 3: Real-World Experiments
The results of this experiment reveal that one or more of our proposed algorithms for our FDR-SVM attain the highest F-1 score for all datasets. However, we observe that ADMM outperforms SM on most datasets, which was not the case in Experiment 1. In fact, the only dataset where SM outperforms ADMM is Parkinson’s, which is a dataset with high class imbalance ($[75\%,25\%]$). This suggests that the performance of the FDR-SVM in terms of F-1 score depends largely on the choice of algorithm. Therefore, in a real-world setting the choice of algorithm should be made based on factors such as guaranteed theoretical convergence, runtime until convergence, and the results of initial preliminary experiments on the dataset under study. We also observe that the performance of ADMM-SC can be significantly lower than that of ADMM on datasets such as BCW and UKM, but it also can be almost identical to it on datasets such as Banknote, Mammographic Mass, and Rice. This highlights that guaranteed theoretical convergence need not always come at a significant sacrifice in model performance, further emphasizing the need for preliminary experimentation when one of those algorithms is being chosen for a real-world problem. Another observation we make is that ADMM-REG attains the same performance as ADMM on some datasets, whereas it falls short of it on most. As demonstrated in Experiment 1, this is most likley due to label uncertainty levels. More specifically, ADMM-REG will perform very similarly to ADMM on datasets where the labels have some uncertainty, whereas it will be outperformed on datasets with noisy labels. Finally, a very surprising observation is that the centralized model does not attain the peak F-1 score on most datasets. This again suggests that our proposed MoWB ambiguity set may lead to performance improvements over the traditional Wasserstein ball in some settings.

### v2

@@ 5.3.2 ADMM Algorithm
This proof follows a very similar strategy to that of Theorem 2. We begin by noting that the strongly convex variant of the local model problem in (12d) equipped with the $\ell_{1}$-norm can be written as a quadratically constrained quadratic problem (QCQP) with $N_{g}+2P+3$ decision variables (including slack variables) and $32N_{g}+2P+$ 3 constraints. When solved via the barrier method equipped with the log barrier function and Newton updates [44], this problem would have the following worst-case time complexity
@@ 6.1.1 Global Parameters
The results show that the SM algorithm outperforms the other distributed models, and is comparable to the central model. The SM algorithm also demonstrates lower sensitivity to variations in $\gamma$ compared to the sensitivity of the ADMM algorithms to changes in $\rho$. This is reflected in its ability to maintain peak performance across a wider range of parameter values. However, this performance comes at the cost of increased training time, as the SM algorithm often requires more iterations than the ADMM algorithms to reach peak performance. Additionally, the SM algorithm appears to be significantly less affected by client and class imbalance, or their different combinations, making it a strong choice in environments where such imbalances are prevalent.
We also observe that the ADMM-trained FDR-SVM model is consistently outperformed by its regularized counterpart (ADMM-REG), except in the noisy labels scenario. This can be explained by the fact that that the DR model accounts for label uncertainty via the label flipping cost parameter $\kappa$, which is lacking in the regularized model [7]. This can cause the FDR-SVM with a relatively low $\kappa$ to be overly conservative when all training labels are correct. In fact, this is further supported by our observation that the FDR-SVM slightly outperforms its regularized counterpart in the noisy label setting. We note, however, that with appropriate local parameter choices ADMM can still outperform ADMM-REG as we show in the local parameter experiment. We also observe that ADMM-SC is consistently outperformed by both ADMM and ADMM-REG. This is likely due to the redundant regularization that is introduced by the extra strongly convex term. Notably, the performance of ADMM-SC declines more sharply than ADMM and ADMM-REG, as $\rho$ increases. This happens because increasing $\rho$ leads to a corresponding increase in $\tau$, which amplifies the effect if the redundant regularization, thus negatively impacting model performance.
Finally, we make the unintuitive observation that ADMM and ADMM-REG slightly outperform the centralized model in the noisy label setting. This unexpected result may suggest that the centralized model is particularly sensitive to its parameters in the noisy label setting. Another plausible explanation is that our proposed FDR-SVM uses a different ambiguity set than the centralized model. This suggests that our proposed ambiguity set is more successful in modeling the uncertainty in the data when different clients have different distributions governing their local data.
@@ 6.1.2 Local Parameters
A key insight from these results is that ADMM is outperformed by ADMM-REG, at low $\kappa$ values. However, as $\kappa$ increases, ADMM matches and in sometimes exceeds the performance of ADMM-REG algorithm. This supports our hypothesis—in the previous Global Parameters experiments—that the reason ADMM-REG outperformed ADMM is due to the over-conservatism of the DR model equipped atlow $\kappa$ values. This result confirms that the FDR-SVM performs just as well as its regularized counterpart, if not better, in many settings including uncertain training labels. We also observe that SM is less sensitive than ADMM to the choice of $\kappa$ for all settings except for the noisy label setting, exhibiting almost opposite behavior to ADMM. Moreover, SM is also relatively highly sensitive to the choice of local Wasserstein radius. Indeed, in most settings the performance of SM improves as the local radius decreases. However, for noisy labels, the performance drops as the local radius decreases. This is likely due to the need for larger local ambiguity sets to account for both feature and label uncertainty. Additionally, we observe again that SM’s peak performance is higher than that of ADMM in all settings except for noisy labels. This is because the results of the global parameter experiment show that SM achieves peak performance in this setting with $\gamma=1\times 10^{1}$. However, in this experiment we test with $\gamma=1\times 10^{2}$. By referencing results from the previous experiments, we observe that SM still performs just as well as ADMM with appropriately chosen parameters. We also observe that ADMM-SC is consistently outperformed by all the other models for most parameter values, which again is likely due to redundant regularization. Interestingly, in some cases, the error bars of the ADMM-SC surpasses all the other models in performance, including the central model. This further suggests that the MoWB ambiguity set can characterize uncertainty more successfully compared to the classical Wasserstein ball in certain settings.
@@ 6.2 Experiment 2: Scalability Experiment
This experiment examines the scalability of the algorithms and models under study using simulated data generated via the make_classification module [52]. The data distribution across clients and classes is identical to that utilized in the nominal setting of the previous experiment. We examine the following settings: a) Increasing clients [fixed training samples]: $N=1000$, $P=4$, $G\in\{10,20,30,40,50\}$; b) Increasing clients [increasing training samples]: $N=100G$, $P=4$, $G\in\{10,20,30,40,50\}$; c) Increasing training samples: $G=10$, $P=4$, $N\in\{1000,1500,2000,2500,3000\}$, and d) Increasing features: $N=4$, $G=10$, $P\in\{4,6,8,10,12\}$.
It is important to note that the subgradient method lacks a practically implementable stopping criterion [60], and similarly, no stopping criterion is provided for multi-block ADMM [50]. To ensure consistency across all algorithms, we evaluate scalability by calculating the mean runtime (in seconds) of each algorithm over $50$ repetitions. For each repetition, we generate new random training and test sets, and measure how long it takes to converge to the model with the highest mCCR. For the SM algorithm, we test performance for $T\in\{140,180,220\}$ and $\gamma\in\{1\times 10^{b}\}_{b=1}^{3}$. For the ADMM algorithms, we test performance for $T\in\{10,20,30\}$ and $\gamma\in\{1\times 10^{b}\}_{b=-3}^{-1}$. Across all algorithms, we fix $\varepsilon_{g}=\frac{1}{10N_{g}}$ and $\kappa=0.25$. The central is tuned in the same way as in the previous experiment, and the runtime that is reported reflects the time taken to solve the optimization problem. The results of this study are reported in Figure 3.
@@ 6.3 Experiment 3: Real-World Experiments
The results of this experiment reveal that one or more of our proposed algorithms for our FDR-SVM attain the highest F-1 score for all datasets. However, we observe that ADMM outperforms SM on most datasets, which was not the case in Experiment 1. In fact, the only dataset where SM outperforms ADMM is Parkinson’s, which is a dataset with high class imbalance ($[75\%,25\%]$). This suggests that the performance of the FDR-SVM in terms of F-1 score depends largely on the choice of algorithm. Therefore, in a real-world setting the choice of algorithm should be made based on factors such as guaranteed theoretical convergence, runtime until convergence, and the results of initial preliminary experiments on the dataset under study. We also observe that the performance of ADMM-SC can be significantly lower than that of ADMM on datasets such as BCW and UKM, but it also can be almost identical to it on datasets such as Banknote, Mammographic Mass, and Rice. This highlights that guaranteed theoretical convergence need not always come at a significant sacrifice in model performance, further emphasizing the need for preliminary experimentation when one of those algorithms is being chosen for a real-world problem. Another observation we make is that ADMM-REG attains the same performance as ADMM on some datasets, whereas it falls short of it on most. As demonstrated in Experiment 1, this is most likley due to label uncertainty levels. More specifically, ADMM-REG will perform very similarly to ADMM on datasets where the labels have some uncertainty, whereas it will be outperformed on datasets with noisy labels. Finally, a very surprising observation is that the centralized model does not attain the peak F-1 score on most datasets. This again suggests that our proposed MoWB ambiguity set can lead to performance improvements over the classical Wasserstein ball in some settings as hypothesized in Remark 1.